OpenStudy (anonymous):

A projectile is fired from vertically upward from the ground level with an initial velocity if 16 ft/sec, how long does it take for the projectile to hit the ground?

OpenStudy (anonymous):

So what is your guess ?

OpenStudy (anonymous):

@TheAni-Trix lemme give you a hint! You have been given u = initial velocity = 16 ft/sec and its obivous that its v = final speed = 0 ft/s and acceleration here would be gravitational acceleration ...i.e. g which is equals to \[9.8 m/s ^{2}\] so now use the equation v = u + at to calculate t and multiply that t with 2 to get your answer..:)

OpenStudy (anonymous):

@TheAni-Trix ... try to solve it :)

OpenStudy (anonymous):

Why do I multiply t by 2?

OpenStudy (anonymous):

@Shizen

mathslover (mathslover):

TheAnt-Trix, you need to multiply "t" by 2 as the ball will first go up (ascent) and then will come down (descent) So, time of complete flight = Time of Descent + Time of Ascent

mathslover (mathslover):

Or Time (T) = 2* t t = Time of Ascent/Descent

OpenStudy (anonymous):

cos the question needs the complete time of the projectile going up and then returning back

OpenStudy (anonymous):

Oh okay, so does that mean that the time to go upward was 1/2?

OpenStudy (anonymous):

it was half the time taken by the projectile for returning back :)